OpenStudy (anonymous):

Factor the algebraic expression below in terms of a single trigonometric function. csc 2x - 1

OpenStudy (raden):

csc(2x) - 1 or csc^2 (x) -1 ??

OpenStudy (anonymous):

csc^2 (x) -1 Sorry about that @raden

OpenStudy (raden):

yea, just do like this : a^2 - b^2 = (a+b)(a-b) does help ?

OpenStudy (raden):

it means a=csc(x) and b=1

OpenStudy (anonymous):

I'm sorry I'm still confused.

OpenStudy (raden):

csc^2 (x) -1 = (cscx +1)(cscx -1)